10Ω 2. Calculate the current I3 for the circuit at left. 120V 200 S0 a. 0.5 A b. 1.0 A c. 2.0 A d. 2.5 A e. 3.0 A
Answer #1

Applying KVL in upper loop
10 I1  + 20I1 -120 + 15 I2 = 0
30I1 + 15 I2 = 120 ---------(1)
Now in the lower loop
15 I2 + 5 I3+ 25I3 - 60 = 0
15I2 + 30I3 = 60 -----------(2)
Now applying KCL
we know I2 = I1 + I3
so 1and 2 becomes
30I1 + 15 (I1 + I3 )= 120
45 I1 + 15I3 = 120 --------(3)
15(I1 + I3 )+ 30I3 = 60
15I1 + 45I3  = 60 -----------(4)
On solving 3 and 4
we get
I1 = 2.5 A and I3 = 0.5 A  
Hence correct option is a